OpenStudy (anonymous):

Funny, and somewhat tough question I found: Suppose that in January 2013, there were 5000 radio stations in the US that played Justin Bieber's music, and that In July 2013, there were only 4000 radio stations in the US that played Justin Bieber's music. Assuming that the rate of decline is proportional to the number of radio stations which play Justin Bieber's music, when would there only be 2000 radio stations which play Justin Bieber's music?

OpenStudy (anonymous):

Options: July 2014 January 2014 July 2015 January 2015

OpenStudy (kc_kennylau):

Let the rate of decline be d

OpenStudy (kc_kennylau):

Between January and July 2013 the rate was 0.8

OpenStudy (kc_kennylau):

And it's proportional to 5000

OpenStudy (kc_kennylau):

So d/5000=constant

OpenStudy (anonymous):

I hope you can figure it out. My answer was not right. I got September 2015.

OpenStudy (kc_kennylau):

=0.00016

OpenStudy (kc_kennylau):

So d=0.00016*b

OpenStudy (kc_kennylau):

July 2013: d=0.00016*4000=0.64, b=4000*0.64=2560

OpenStudy (kc_kennylau):

January 2014: d=0.00016*2560=0.4096, b=2560*0.4096=1048.576

OpenStudy (kc_kennylau):

Wait...

OpenStudy (kc_kennylau):

The decline was 1000

OpenStudy (anonymous):

My answer is C before any other answers are thrown out there.

OpenStudy (kc_kennylau):

The rate of decline is proportional... so the rate of decline isn't a constant

OpenStudy (kc_kennylau):

The rate of decline was 0.2, 0.2/5000=0.00004

OpenStudy (kc_kennylau):

So d=0.00004b

OpenStudy (kc_kennylau):

July 2013: d=0.00004*4000=0.16, b=4000*(1-0.16)=3360

OpenStudy (kc_kennylau):

January 2014: d=0.00004*3360=0.1344, b=3360*(1-0.1344)=2908.416

OpenStudy (kc_kennylau):

July 2014: d=0.00004*2908.416=0.11633664, b=2908.416*(1-0.11633664)=2570.06065

OpenStudy (kc_kennylau):

January 2015: d=0.00004*2570=0.1028, b=2570*(1-0.1028)=2305.804

OpenStudy (kc_kennylau):

July 2015: d=0.00004*2305.804=0.09223216, b=2305.804*(1-0.09223216)=2093.13472

OpenStudy (kc_kennylau):

Close enough; July 2015

OpenStudy (anonymous):

What do you mean close enough.

ganeshie8 (ganeshie8):

i smell a diff eqn..

OpenStudy (kc_kennylau):

\[\Large b_{n+1}=(1-0.00004b_n)b_n\]

OpenStudy (kc_kennylau):

Write an explicit formula for that, b_0=5000

OpenStudy (anonymous):

January 2015